OpenStudy (anonymous):

Solve the system of linear equations below. 2x+3y=2 x+6y=4

OpenStudy (anonymous):

Okay, let's use elimination method.

OpenStudy (anonymous):

2(2x + 3y = 2) 4x + 6y = 4 - x + 6y = 4 -3x = 0 No real solutions.

OpenStudy (anonymous):

Error, does not compute. Shutting down.

OpenStudy (anonymous):

*dies*

OpenStudy (anonymous):

Lemme try sub method.

OpenStudy (anonymous):

ok now what

OpenStudy (anonymous):

There is one answer that has a 0

OpenStudy (anonymous):

2x+3y=2 2x = 2 - 3y x = 1 - 3/2y 2(1 - 3/2y) + 3y = 2 1 - 3y + 3y = 2 1 = 2 Error, does not compute.

OpenStudy (anonymous):

No, every value of x = 0 for the other one. So no solutions.

OpenStudy (anonymous):

No solutions for this one.

OpenStudy (anonymous):

There isn't a no solution answer :c

OpenStudy (anonymous):

Is there an infinite solution?

OpenStudy (anonymous):

What are the choices?

OpenStudy (anonymous):

A) x=0, y=2/3 B) x=2, y=-2/3 C) x=4/3, y=2/9 D) x=4, y=-2

OpenStudy (anonymous):

Okay, since I'm not getting an answer, I'll sub in those values. 2x + 3y = 2 0 + 6 = 2 6 = 2 FALSE 2x + 3y = 2 4 - 2 = 2 2 = 2 TRUE Double Check: x + 6y = 4 2 + 6(-2/3) = 4 2 - 4 = 4 -2 = 4 FALSE Double Check FAILED 2x + 3y = 2 2*4/3 + 3*2/9 = 2 FALSE 2x + 3y = 2 8 - 6 = 2 2 = 2 TRUE Double Check: x + 6y = 4 4 - 12 = 4 -8 = 4 FALSE Double Check FAILED

OpenStudy (anonymous):

So, no solution. I don't know what to tell you. Contact your teacher for help and see if they have a different answer.

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (anonymous):

No problem, I'm sorry I can't see any other answers.

OpenStudy (anonymous):

I'm going to run everything through three separate online calculators and see what they think.

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

Okay, wait, I see an error that I did.

OpenStudy (anonymous):

2x + 3y = 2 0 + 2 = 2 True Double check: x+6y=4 0 + 6(2/3) = 4 True

OpenStudy (anonymous):

My mistake was: 3 * 2/3 = 6 3 * 2/3 is 2, because 2 / 3 * 3 is 2.

OpenStudy (anonymous):

Sorry about that. :)

OpenStudy (anonymous):

A is correct.

OpenStudy (anonymous):

It's okay :D thanks!